OpenStudy (anonymous):

Determine if the side lengths 6, 11, and 19 form a triangle. If it is a triangle, classify it by its sides.

OpenStudy (anonymous):

take sum of squares of two smaller sides and see whether they are equal or greater or lesser than the square of lager side. like in this case 6^2+11^2=36+121=157 and 19^2=361 it means that the sum of squares of two smaller sides is less than square of larger side i.e \[157<361\] if the sum of the squares of the two smaller sides is less than the square on the largest side, you will have an "obtuse triangle"

OpenStudy (anonymous):

@beezz12 hope you get it